the ceiling of an arena is 45m above the floor. what is the minimum speed that a thrown ball would need to just reach the ceiling
1
Expert's answer
2020-02-05T12:26:40-0500
"v^2=2gh"

"v^2=2(9.8)(45)"

"v=29.7\frac{m}{s}"